OpenStudy (samsan9):

A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 130 feet per second. What is the maximum height, in feet, the ball will attain?

OpenStudy (tanya123):

h(t) = - 16t² + v0 + h0 where h = height, t = time, v0 = initial velocity and h0 = initial height v0 = 130 h0 = 2 h(t) = - 16t² + 130t + 2 Max. Ht.: h'(t) = 0: h'(t) = - 32t + 130 - 32t + 130 = 0 - 32t = - 130 t = - 130 / - 32 t = 4.06 sec.

OpenStudy (samsan9):

you are awesome :D

OpenStudy (tanya123):

Thank you, so are you!